3.00 points Tracie invested $14,000 five years ago in TLK stock. How much is her investment worth today if she has earned a 14.2 percent rate of return? O $23,811.80 O $23,940.00 $31,054.51 O $27,193.09
value: 3.00 points Sue invested $5,100 at 7 percent interest for 4 years. Phil invested $5,100 at 7 percent interest for 8 years. How much more interest will Phil earn than Sue? O $1,428.00 O $2,07769 O $1,109.02 $1,723.29
value: 3.00 points You want to invest $13,000 for 5 years. How much additional interest will you earn if you invest at 6 percent compound interest rather than 6 percent simple interest? O $506.52 O $496.93 $780.00 $0
Answer #1

Answers:

  1. Formula for calculating present worth of an investment:

PV = X ((1+r) ^n)

= 14000(1+0.142) ^5)

= 14000*1.9424

= 27193.09, option d is the correct answer.

  1. Formula for calculating interest:

I = p*r*t

For sue:

I = 5100*0.07*4 = 1428

For Phil:

I = 5100*0.07*8 = 2856

Hence interest earned by Phil = 2856 – 1428 = 1428, option a is the correct answer.

  1. Simple interest:

= 13000*5*0.06 = 3900

Compound interest:

C = (p (1 +r) ^n – 1)

= 13000(1+0.06) ^5 – 1)

= 4396.93

Difference = 4396.93 – 3900 = 496.93, option b is the correct answer.